differentiate the following function with respect to x
tan^-1(a+x/1-ax)

Grade:12

2 Answers

Rinkoo Gupta
9 years ago

y=tan-1{(a+x)/(1-ax)}

dy/dx=1/[1+{(a+x)/(1-ax)}2].[(1-ax).1-(a+x).(-a)]/(1-ax)2

=[1/{(1-ax)2+(a+x)2}].[1-ax+a2+ax]

=(1+a2)/(1+a2x2+a2+x2)

=(1+a2)/(1+a2)(1+x2)

=1/(1+x2)

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty

Arun Kumar IIT Delhi
9 years ago
Hi

Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty